OpenStudy (anonymous):

1. Using the following system of equations : 4x +2x = 6 2x + y= 3 A: Find the solution(s) algebraically showing all work (4pts) B: State whether it is independent consistent, dependent consistent, or inconsistent. (2pts) C: Demonstrate how to check the systems of solutions, showing all work. (4pts) A: Find the solution(s) algebraically showing all work (4pts) B: State whether it is independent consistent, dependent consistent, or inconsistent. (2pts) C: Demonstrate how to check the systems of solutions, showing all work. (4pts)

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

For A, you can simplify the first equation by combining like terms. It'll look like: 4x+2x=6 6x=6 x=1 Then you can plug x into the second equation. 2(1)+y=3 y=1 So the solution to the system of equations is (1,1).

OpenStudy (anonymous):

For B, if lines are parallel (they won't ever intersect) they are inconsistent. Sometimes you'll be given the same line twice, that means the system is dependent. And it's independent if they're two different lines that intersect. So in this case, they are independent.

OpenStudy (anonymous):

thank you what about c

OpenStudy (anonymous):

For C, you can check it by plugging your x and y values [from the ordered pair (x,y)] into one of the equations that has both an x and a y. So for this, we'd use the second equation. 2x+y=3 2(1)+1=3 2+1=3 3=3, since that is true, that means your systems of solutions is correct. Hope all that helps, lemme know if you have more questions! :)

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

No prob!

OpenStudy (anonymous):

@bruhhh could yo help with another